OpenStudy (anonymous):

If a tennis ball has a diameter of approximately 6.5 centimeters, what is its surface area? 530.93 square centimeters 265.46 square centimeters 132.73 square centimeters 862.76 square centimeters

OpenStudy (dumbcow):

SA = 4pi*r^2

OpenStudy (anonymous):

Um, I dont even know how to apply this equation but thanks

OpenStudy (dumbcow):

then how are you going to figure out the surface area if you don't know how to use the formula ? all you have to do is plug in the radis of the ball

OpenStudy (anonymous):

Well if I knew that, would I really be here. Thanks! I'll skip the problem.

OpenStudy (dumbcow):

haha ok radius is half the diameter --> 1/2*6.5 = 3.25 SA = 4*pi*(3.25)^2 pi is 3.14 SA = 4(3.14)(3.25)^2 = 132.7 For surface area and volume you need to use and be familar with the formulas, thats all i was saying

OpenStudy (anonymous):

So much is easier! I didn't know how to find the radius so I didn't know to plug in the numbers. I just wasn't familiar with the equations. Thanks :)